jerome mixes a solution that is 8% alcohol with a solution that is 14% alcohol to make 600 milliliters of a solution that is 12% alcohol. How many milliliters of each solution does he mix together?
Answer by lwsshak3(11628)

jerome mixes a solution that is 8% alcohol with a solution that is 14% alcohol to make 600 milliliters of a solution that is 12% alcohol. How many milliliters of each solution does he mix together?
***
let x=amt of 8% solution to mix
600-x=amt of 14% solution to mix
..
8%x+14%(600-x)=12%*600
.08x+84-.14x=72
.06x=12
x=200
600-x=400
amt of 8% solution to mix=200 ml
amt of 14% solution to mix=400 ml